OpenStudy (anonymous):

Phillip received 75 points on a project for school. He can make changes and receive two-tenths of the missing points back. He can make corrections as many times as he wants. Create the formula for the sum of this geometric series and explain your steps in solving for the maximum grade Phillip can receive. Identify this as converging or diverging.

OpenStudy (dumbcow):

hmm do we know how successful his corrections are each time around? for example if he was perfect on corrections first time around...that would add 25*0.2 = 5 points for total of 80 points

OpenStudy (anonymous):

mmm well it says maximum grade so i'm guessing after corrections they were all correct

OpenStudy (dumbcow):

right so max is 80 but not sure how to set up geometric series since what it being multiplied by 0.2 varies "He can make corrections as many times as he wants"

OpenStudy (dumbcow):

here are general formulas \[P = a(1+.2 +.2^{2}+....2^{n})\] \[Sum = \frac{a(1-.2^{n})}{1-.2}\]

OpenStudy (anonymous):

so would it be 80=75(1-.2^n)/1-/2?

OpenStudy (dumbcow):

no that wont work , sorry im not sure what they want here as far as geometric sequence i just gave the general form of what it should look like and logically a max of 80 makes sense sorry i cant help further

OpenStudy (anonymous):

Okay, well thank you for trying (: